250 mg of caco3 was dissolved in hcl and the solution was made 250 ml with distilled water. 50 ml of the above solution required 20 ml of edta solution [standardization part]. 50 ml of hard water sample consumed 25 ml of edta with ebt as indicator [total hardness]. calculate the total hardness of water sample in ppm.
1
Expert's answer
2020-12-15T07:51:51-0500

1. Standard calcium solution.

(0.250 g CaCO3) (1 mol CaCO3 / 100.09 g CaCO3)(1 mol Ca2+ / 1 mol CaCO3) = 0.00250 mol Ca2+ ions

Molarity of Ca2+ = 0.00250 mol Ca2+ / 0.250 L = 0.01 M Ca2+ ions

2. Standardization of EDTA.

(0.01 mol Ca2+ / L)(0.050 L / 1)(1 mol EDTA / 1 mol Ca2+) = 0.0005 mol EDTA

Molarity of EDTA = 0.0005 mol EDTA / 0.020 L = 0.025 M EDTA

3. Ca2+ concentration in an unknown solution CaCO3 reported as ppm CaCO3.

(0.025 mol EDTA / L)(0.025 L / 1)(1 mol Ca2+ / 1 mol EDTA) = 0.000625 mol Ca2+

0.000625 mol Ca2+ (1 mol CaCO3 / 1 mol Ca2+)(100.09 g CaCO3 / 1 mol CaCO3) = 0.062556 g CaCO3

ppm CaCO3 = mg CaCO3 / L Solution = 0.062556 g CaCO3 (1000 mg CaCO3 / 1.00 g CaCO3) (1 / 0.050 L Solution) = 1251

1251 ppm CaCO3